Billy is 6 years younger than Amy. The sum of their ages is 22. How old will each of them be in 8 years?
Answer by jim_thompson5910(35256)

B = A - 6

A + B = 22

A + A - 6 = 22

2A = 28

A = 14

B = A - 6

B = 8

Billy: 8 years old (currently)

Amy: 14 years old (currently)

In eight years, Billy will be 16 yrs old and Amy will be 22 years old.